You dissolve 1.0 mole of a substance in water to a total volume of 1,000 ml. The concentration of the solutions is ______. (Chose all that apply).

1 M

0.1 mole in 100 ml

10 mmole in 10ml

1 mmole in 1 ml
Answer #1

You dissolve 1.0 mole of a substance in water to a total volume of 1,000 ml. The concentration of the solutions is (all answers are correct)

1 M (1 M substance in 1 liter = 1 molar solution)

0.1 mole in 100 ml (0.1M = 1/10thof 1mole and 100ml = 1/10th of 1liter ;final concentration same)

10 mmole in 10ml (10mM= 1/100th of 1mole and 10ml = 1/100th of 1 liter ; final concentration same)

1 mmole in 1 ml (1mM= 1/1000th of 1 mole and 1ml = 1/1000th of 1 liter ; final concentration same)